# Xue Liu, , Zhu Han, , and Charlie Zhang

### v1

@@ Abstract
Generative artificial intelligence (GAI) is a promising technique towards 6G networks, and generative foundation models such as large language models (LLMs) have attracted considerable interest from academia and telecom industry. This work considers a novel edge-cloud deployment of foundation models in 6G networks. Specifically, it aims to minimize the service delay of foundation models by radio resource allocation and task offloading, i.e., offloading diverse content generation tasks to proper LLMs at the network edge or cloud. In particular, we first introduce the communication system model, i.e., allocating radio resources and calculating link capacity to support generated content transmission, and then we present the LLM inference model to calculate the delay of content generation. After that, we propose a novel in-context learning method to optimize the task offloading decisions. It utilizes LLM’s inference capabilities, and avoids the difficulty of dedicated model training or fine-tuning as in conventional machine learning algorithms. Finally, the simulations demonstrate that the proposed edge-cloud deployment and in-context learning task offloading method can achieve satisfactory generation service quality without dedicated model training or fine-tuning.
@@ I Introduction
Generative AI (GAI) has received considerable attention recently, which is capable of analyzing complex data distributions and generating similar new content. Due to its promising features, existing studies have started exploring GAI-enabled 6G networks, e.g., pre-trained generative models for semantic communication [1], and generative neural networks for air-to-ground channel modelling [2]. As a sub-field of GAI, generative foundation models, especially large language models (LLMs), have attracted interest from both academia and telecom industry. On the academia side, foundation models and LLMs are used for network intrusion detection [3], reconfigurable intelligent surface [4], and integrated satellite-terrestrial networks [5]. On the industry side, telecom companies have started applying foundation models. For example, Apple will bring ChatGPT to iPhones with OpenAI, and Qualcomm has developed a new mobile platform to support LLMs.
@@
The above progress of academia and telecom industry has demonstrated the great potential of GAI foundation models such as LLMs in 6G networks [1, 2, 3, 4, 5]. However, despite the advancement, some fundamental and crucial problems are still not investigated, e.g., practical deployment of GAI foundation models within 6G network architecture, and evaluating the service delay of these generation services in wireless environments. In particular, large GAI models such as GPT-4 usually have billions of parameters, requiring considerable storage and computational resources. The practical deployment of these large GAI models is critical to support various applications in wireless networks, i.e., network intrusion detection, generation, and management tasks in [3, 4, 5]. On the other hand, mobile users may send various service requests over wireless channels with diverse preferences, e.g., question-answering tasks need higher accuracy and chatting tasks expect lower delay. Therefore, using one GAI model, i.e., a single LLM, to service the requests of all mobile users is impractical, leading to lower service quality and efficiency.
@@
To this end, we consider a novel edge-cloud collaboration deployment strategy. We assume small-scale LLMs such as Llama3-8B are deployed at network edge servers of base stations (BSs), aiming to process tasks efficiently with lower delay. By contrast, large-scale LLMs, e.g., Llama3-70B and GPT-4, are deployed in the central cloud with abundant computational resources. These large-scale LLMs can generate high-quality content for quality-preferred tasks, but the processing and generation time may be longer [6].
@@
Such an edge-cloud collaboration enables flexible content generation in wireless networks, but it also involves task offloading decisions, i.e., generating content at network edge or offloading tasks to central cloud. Inspired by the recent progress of LLM-based optimization [7], this work further explored in-context learning-based decision-making. Specifically, it uses LLMs to learn from formatted natural language demonstrations and improve the performance on target tasks[8]. Compared with existing machine learning (ML) methods, in-context learning has several advantages: a) Avoiding the complexity of model training and fine-tuning, a well-known bottleneck of conventional ML techniques; b) Following human language instructions to formulate and solve problems, which is far beyond the capabilities of other ML algorithms.
@@
1) Firstly, to the best of our knowledge, this work is the first to model the service delay of foundation GAI models in wireless networks, including the communication models for content transmission, and LLM inference models for content generation. It provides a specified metric to evaluate the delay experienced by mobile users. 2) Secondly, we propose a novel in-context learning method for generation task offloading. It avoids the complexity of dedicated model training and fine-tuning in existing ML techniques, using natural language for network management. Finally, the simulations demonstrate that the proposed edge-cloud deployment can successfully handle various generation tasks from mobile users, and the proposed in-context learning-based method can achieve satisfactory service quality.
@@ II System Model
Fig. 1 presents the proposed system model, in which mobile users can require various LLM services over wireless networks. Small-scale LLMs are deployed at the network edge to process tasks efficiently with lower delay. By contrast, large-scale LLMs are deployed in the central cloud to handle tasks with high-quality requirements. Therefore, the edge servers must make offloading decisions properly, i.e., processing generation tasks locally or offloading them to the central cloud.
@@ II-A Communication Model for Content Transmission
where $n^{token}_{k,i}$ is the LLM output token numbers for input prompt $i$, $C_{j,k}$ is the link capacity for download transmission between user $k$ and BS $j$, $s^{token}$ is the byte size per token[9]. $\frac{n^{token}_{k,i}s^{token}}{C_{j,k}}$ represents the transmission delay from BSs to users. $t_{back}$ is the backhaul delay caused by task offloading. $\alpha_{k,i}$ is the task offloading decision: $\alpha_{k,i}=1$ means offloading tasks to central cloud, while $\alpha_{k,i}=0$ indicates network edge implementation. We assume a fixed $t_{back}$ in (1), which is a setting used in many task offloading-related studies[10]. The link capacities $C_{j,k}$ can be calculated by[11]
@@ II-C Problem Formulation
Finally, (4) includes both radio resource allocation and task offloading decisions, in which $z_{j,q,k}$ for radio resource allocation and $\alpha_{k,i}$ for task offloading. For radio resource allocation, we apply a classic proportional fairness algorithm because: 1) proportional fairness is a practical method that has been widely used in many existing studies [11]; 2) this work aims to understand generative foundation model service properties, and it is reasonable to apply a well-known resource allocation method to better focus on foundation models. For task offloading, we propose an in-context learning-based task offloading approach in the following Section III.
@@ III-A Prompting System Design
Task description for LLM-enabled task offloading Task goal: You are given a decision-making task for computational task offloading. You need to select between two decisions: “local” or “offload”. Task definition: You have to consider the condition of each case, including the following keywords: Keyword 1: Task types, Keyword 2: Estimated output token size. Rules: Now I give you a new condition to solve, please reply “local” or “offload” only without other words.
@@
Step 2: Example design. The task description $D$ will be combined with the example set =$\mathcal{E}=\{E_{1},E_{2},...,E_{|\mathcal{E}|}\}$ as a meta prompt input to the LLM agent, producing task offloading decisions $\alpha$. The example $E_{e}\in\mathcal{E}$ is defined by
@@
where $T^{Target}$ is the target delay, and $t^{total}$ is the objective function defined in (4) and $t^{total}=t^{tran}_{k,i}-\alpha_{k,i}t^{cloud}_{k,i}-(1-\alpha_{k,i})t^{edge}_{k,i}$. Note that we consider a penalty term $r^{penalty}$ to (10) if the constraints in the problem formulation in (4) are violated, which is a widely used approach in reinforcement learning studies to handle optimization constraints[11]. Therefore, (10) provides a comprehensive metric to minimize the total delay under the constraints.
Step 3: Experience evaluation and reply. Given the offloading decision, the current network operation results become a new experience example $E_{current}$, which will be sent to the evaluator module as in Fig. 3. Specifically, $E_{current}$ is considered a bad decision if the constraints in (4) are not satisfied; otherwise, $E_{current}$ is sent to the prioritized experience replay module for further evaluation, which will be introduced in the following Subsection III-B.
Step 4: Meta prompt updating. The experience pool in Step 3 will serve as a new example set $\mathcal{E}_{next}$, and the LLM agent will generate the next decision using the same task description $D$ and the updated example set $\mathcal{E}_{next}$. Steps 3 and 4 will be repeated until generating satisfied performance.
@@ III-B Prioritized Experience Replay and Exploration Strategies
The above two strategies guarantee that the experience pool only includes examples with the best performance, providing the most useful demonstrations to LLM agents. Additionally, exploration and exploitation are fundamental problems in many optimization tasks. In this work, the prioritized experience replay requires better examples to improve the experience pool. Here, we apply an epsilon-greedy method to balance the exploration and exploitation: actions are randomly selected with probability $\epsilon$; otherwise, LLM will make decisions. Such a policy can constantly explore new actions to produce better examples and, meanwhile, make decisions based on existing LLM experience to maximize the reward. In summary, the prioritized experience replay and epsilon-greedy method can send the best examples found to LLMs, and also explore new examples to improve the experience pool.
III-C Baseline Algorithm
This subsection presents 3 baseline algorithms. Baseline 1: Latest experience-based in-context learning, using the latest experience as examples in the prompt. Baseline 2: In-context learning without exploration, and all decisions are made by LLMs. Baseline 3: We consider deep reinforcement learning (DRL) as an optimal baseline since DRL techniques have been very widely applied to solve various network optimization problems, in which the state is defined by (8), the action is offloading decision, and the reward is shown as (10).
@@ IV Performance Evaluation
(c) Comparisons under different experience replay methods using Llama3-8B.
(d) Comparisons under different exploration strategies using Llama3-8B.
(e) Average service delay with different prompt size using different LLM combinations
(f) Average service delay with different tasks proportions using different LLM combinations
@@ IV-A Simulation Settings
We assume 20 users are randomly distributed in a cell based on 3GPP urban networks. The mobile users have two kinds of generation tasks: regular tasks such as chatting and translation, and quality-preferred tasks such as reasoning and knowledge-related services [^2] Such a classification aligns well with the evaluation methods for LLMs, e.g., Chatbot Arena shows the general capabilities, and Multi-task Language Understanding (MMLU) assesses the reasoning and knowledge capabilities.. The average output size is 1,000 tokens, and each token corresponds to about 4 bytes according to an OpenAI report [9]. We evaluate 2 LLMs for decision-making: Llama3-8B as a small-scale model that can be deployed at the network edge and GPT-4o as the latest large-scale LLM model as a benchmark for LLM-enabled methods. For the network settings, we assume the small-scale edge LLM’s TTFT and TPOT values are 0.23 and 1/75 second, while the values for the cloud large-scale LLM are 0.42 and 1/32. These values are extracted from the long-term statistical performance in [6].
@@ IV-B Simulation Results
We further compare the experience replay and exploration methods. Fig. 4(c) shows that the proposed prioritized experience replay method can obtain a higher reward than using the latest experience. This is because prioritized experience replay can always present the best examples found so far, while the latest experience cannot represent the historical exploration record. Meanwhile, Fig. 4(c) shows the importance of random explorations. epsilon-greedy exploration can explore the environments by trying different decisions, contributing to the collection of good examples. Then, these good examples will be used in prioritized experience replay to guide the future decisions of LLMs. By contrast, without proper exploration methods, the LLM agent may stick with one single action.
Finally, we investigate the LLM service delay under different edge and cloud LLM combinations, indicating various TTFT and TPOT values. The smaller LLMs are deployed at the network edge, and large LLMs are implemented in the cloud. Fig. 4(e) and 4(f) reveal that Gemma 7B + Gemini 1.5 Pro can achieve the lowest overall delay than other combinations. This is because of the low TTFT values of these two LLMs, which are 1/155 second for Gemma 7B and 1/58 second for Gemini 1.5 Pro. Other LLMs have much higher values, i.e., 1/89 second for Llama2-7B and 1/40 second for Llama2-70B.
@@ V Conclusion
GAI is a promising technique for future 6G networks, and this work investigates the generation task offloading problems in 6G edge-cloud. This work first defines the wireless communication system model and the LLM inference model. It also proposes a novel in-context learning method for generation task offloading between network edge and cloud. The simulations demonstrated that edge-cloud collaboration architecture can well satisfy user requirements, and the proposed in-context learning method can achieve satisfactory performance without dedicated model training or fine-tuning.

### v2

# Xue Liu, , Zhu Han, , and Jianzhong (Charlie) Zhang

@@ Abstract
Generative artificial intelligence (GAI) is a promising technique towards 6G networks, and generative foundation models such as large language models (LLMs) have attracted considerable interest from academia and industry. This work considers a novel edge-cloud deployment of foundation models in 6G networks. Specifically, it aims to minimize the service delay of foundation models by radio resource allocation and task offloading, i.e., offloading diverse content generation tasks to proper LLMs at the network edge or cloud. In particular, we first introduce the communication system model, i.e., allocating radio resources and calculating link capacity to support generated content transmission, and then we present the LLM inference model to calculate the delay of content generation. After that, we propose a novel in-context learning method to optimize the task offloading decisions. It utilizes LLM’s inference capabilities, and avoids the difficulty of dedicated model training or fine-tuning as in conventional machine learning algorithms. Finally, the simulations demonstrate that the proposed edge-cloud deployment and in-context learning method can achieve satisfactory generation service quality without dedicated model training.
@@ I Introduction
Generative AI (GAI) has received considerable attention recently, which is capable of analyzing complex data distributions and generating similar new content. Due to its promising features, existing studies have started exploring GAI-enabled 6G networks, e.g., GAI for semantic communication [1] and air-to-ground channel modelling [2]. As a sub-field of GAI, generative foundation models, especially large language models (LLMs), have attracted interest from both academia and industry. For instance, foundation models have been used for network intrusion detection in [3]. Meanwhile, telecom companies have started applying foundation models, e.g., Apple will bring ChatGPT to iPhones with OpenAI, and Qualcomm has developed a mobile platform to support LLMs.
@@
The above progress of academia and telecom industry has demonstrated the great potential of GAI foundation models such as LLMs in 6G networks [1, 2, 3, 4, 5]. However, despite the advancement, some fundamental and crucial problems are still not investigated, e.g., practical deployment of GAI foundation models within 6G network architecture, and evaluating the service delay of these generation services in wireless environments. In particular, large GAI models such as GPT-4 usually have billions of parameters, and practical deployment of these large GAI models is critical to support various applications in wireless networks, i.e., network intrusion detection, generation, and management tasks in [3, 4, 5]. On the other hand, mobile users may send various service requests over wireless channels with diverse preferences, e.g., question-answering tasks need higher accuracy and chatting tasks expect lower delay. Therefore, using one GAI model, i.e., a single LLM, to service the requests of all mobile users is impractical, leading to lower service quality and efficiency.
@@
To this end, we proposed a novel edge-cloud collaboration deployment strategy. We consider small-scale LLMs such as Llama3-8B to be deployed at network edge servers of base stations (BSs), aiming to process tasks efficiently with lower delay, e.g., chatting, information extraction, and content summarization. By contrast, large-scale LLMs, e.g., Llama3-70B and GPT-4, are deployed in the central cloud with abundant computational resources. These large-scale LLMs can generate high-quality content for quality-preferred tasks, such as scientific knowledge, math, and coding-related tasks.
@@
Such an edge-cloud collaboration enables flexible content generation in wireless networks, but it also involves task offloading decisions, i.e., generating content at network edge or offloading tasks to central cloud. Inspired by the recent progress of LLM-based optimization [6], this work further explored in-context learning-based decision-making. Specifically, it uses LLMs to learn from formatted natural language demonstrations and improve the performance on target tasks[7]. Compared with existing machine learning (ML) methods, in-context learning has several advantages: a) Avoiding the complexity of model training and fine-tuning, a well-known bottleneck of conventional ML techniques; b) Following human language instructions to formulate and solve problems, which is far beyond the capabilities of other ML algorithms. In-context learning has been explored to address detection, optimization, prediction tasks, etc [6].
@@
1) Firstly, to the best of our knowledge, this work is the first to model the service delay of foundation GAI models in wireless networks, including the communication models for content transmission, and LLM inference models for content generation. It provides a specified metric to evaluate the delay experienced by mobile users. 2) Secondly, we propose a novel in-context learning method for generation task offloading. It avoids the complexity of dedicated model training and fine-tuning as existing ML techniques, using natural language for network management. Finally, the simulations demonstrate that the proposed edge-cloud deployment strategy and in-context learning-based method can achieve satisfactory generation service quality for mobile users.
@@ II System Model
Fig. 1 presents the proposed system model, in which mobile users can require various LLM services over wireless networks, e.g., chatting, information extraction and summarization, image generation, etc. Note that these tasks may have different requirements for generated content qualities. Small-scale LLMs are deployed at the network edge to process tasks efficiently with lower delay. By contrast, large-scale LLMs are deployed in the central cloud to handle tasks with high-quality requirements. Therefore, the edge servers must make offloading decisions properly, i.e., processing generation tasks locally or offloading them to the central cloud.
@@ II-A Communication Model for Content Transmission
where $n^{token}_{k,i}$ is the LLM output token numbers for input prompt $i$, $C_{j,k}$ is the link capacity for download transmission between user $k$ and BS $j$, $s^{token}$ is the byte size per token[8]. $\frac{n^{token}_{k,i}s^{token}}{C_{j,k}}$ represents the transmission delay from BSs to users. $t_{back}$ is the backhaul delay caused by task offloading. $\alpha_{k,i}$ is the task offloading decision: $\alpha_{k,i}=1$ means offloading tasks to central cloud, while $\alpha_{k,i}=0$ indicates network edge implementation. We assume a fixed $t_{back}$ in (1), which is a setting used in many task offloading-related studies[9]. The link capacities $C_{j,k}$ can be calculated by
@@ II-C Problem Formulation
Finally, (4) includes both radio resource allocation and task offloading decisions, in which $z_{j,q,k}$ for radio resource allocation and $\alpha_{k,i}$ for task offloading. For radio resource allocation, we apply a classic proportional fairness algorithm because: 1) proportional fairness is a practical method that has been widely used in many existing studies; 2) this work aims to understand generative foundation model service properties, and it is reasonable to apply a well-known resource allocation method to better focus on foundation models. For task offloading, we propose an in-context learning-based task offloading approach in the following Section III.
@@ III-A Prompting System Design
Step 2: Example design. The task description $D$ will be combined with the example set =$\mathcal{E}=\{E_{1},E_{2},...,E_{|\mathcal{E}|}\}$ as a meta prompt input to the LLM, producing task offloading decisions $\alpha$. The example $E_{e}\in\mathcal{E}$ is defined by
@@
where $T^{Target}$ is the target delay, and $t^{total}$ is the objective function defined in (4) and $t^{total}=t^{tran}_{k,i}-\alpha_{k,i}t^{cloud}_{k,i}-(1-\alpha_{k,i})t^{edge}_{k,i}$. Here $r^{penalty}$ is a penalty item, aiming to balance the service delay and generated content quality. $r^{penalty}$ is a preset positive value if the constraint in (4) is violated. A high $r^{penalty}$ value means strict requirements on generated content quality, while a lower $r^{penalty}$ value allows lower generation quality to achieve lower service delay. Otherwise, if all constraints are satisfied in (4), $r^{penalty}=0$. Such a definition is also a widely used approach in reinforcement learning studies to handle optimization constraints. Therefore, (10) provides a comprehensive metric to minimize the total delay under the constraints.
Step 3: Experience evaluation and replay. Given the offloading decision, the current network operation results become a new experience example $E_{current}$, which will be sent to the evaluator module as in Fig. 3. Specifically, $E_{current}$ is considered a bad decision if the constraints in (4) are not satisfied; otherwise, $E_{current}$ is sent to the prioritized experience replay module for further evaluation, which will be introduced in the following Subsection III-B.
Step 4: Meta prompt updating. The experience pool in Step 3 will serve as a new example set $\mathcal{E}_{next}$, and the LLM will generate the next decision using the same task description $D$ and the updated example set $\mathcal{E}_{next}$. Steps 3 and 4 will be repeated until algorithm performance converges, which will be analyzed in the following Fig. 4.
@@ III-B Prioritized Experience Replay and Exploration Strategies
Figure 4: The overall optimization procedure of the proposed technique
Fig. 4 summarizes the overall procedure of the proposed techniques. We apply the well-known epsilon-greedy policy to balance exploration and exploitation, which is a fundamental problem in many optimization tasks. Specifically, actions are randomly selected with probability $\epsilon$; otherwise, LLM will make decisions. Combining the epsilon-greedy policy with the proposed experience replay technique can send the best examples found to LLMs, and also explore new examples to improve the experience pool. Therefore, with plenty of explorations, the examples in the experience pool will be constantly improved, and the LLM output will converge if no better examples can be found [^2] Similar assumptions can be found in the convergence analyses of classic tabular-based Q-learning, e.g., “Q-learning is guaranteed to converge when visiting each state-action tuple infinitely times”..
Finally, we present 3 baseline algorithms. Baseline 1: Latest experience-based in-context learning, using the latest experience as examples in the prompt. Baseline 2: In-context learning without exploration, and all decisions are made by LLMs. Baseline 3: We consider deep reinforcement learning (DRL) as an optimal baseline since DRL techniques have been very widely applied to solve various network optimization problems, in which the state is defined by (8), the action is offloading decision, and the reward is shown as (10).
@@ IV Performance Evaluation
(c) Comparisons under different experience replay and exploration methods using Llama3-8B.
(d) Average service delay with different user numbers using different LLM combinations.
(e) Average service delay with different prompt size using different LLM combinations.
(f) Average service delay with different tasks proportions using different LLM combinations.
@@ IV-A Simulation Settings
We assume 20 users are randomly distributed in a cell based on 3GPP urban networks. The mobile users have two kinds of generation tasks: regular tasks such as chatting, translation, and summarization, and quality-preferred tasks such as scientific knowledge reasoning and coding. Such quality-based task classification aligns well with the evaluation schemes for LLMs, e.g., Chatbot Arena and Multi-task Language Understanding (MMLU). The average output size is 1,000 tokens, and each token corresponds to about 4 bytes according to an OpenAI report [8]. We evaluate 2 LLMs for decision-making: Llama3-8B as a small-scale model that can be deployed at the network edge and GPT-4o as the latest large-scale LLM model as a benchmark for LLM-enabled methods. For the network settings, we assume the edge LLM’s TTFT and TPOT values are 0.23 and 1/75 second, while the values for the cloud LLM are 0.42 and 1/32 [11].
@@ IV-B Simulation Results
We further compare the experience replay and exploration methods. in Fig. 5(c). It shows that the proposed prioritized experience replay method can obtain comparable performance as optimal DRL baseline values. Specifically, epsilon-greedy exploration can try different decisions and collect good examples. These good examples will then be used in prioritized experience replay to guide the future decisions of LLMs. By contrast, latest experience reply and no-exploration methods present lower rewards as shown in the lower part of Fig. 5(c).
Finally, Figs. 5(d), 5(e) and 5(f) present the service delay of various LLMs under different user numbers, size of output tokens, and task distributions. In particular, when the proportion of quality-preferred tasks increases, these generation tasks must be offloaded to large-scale LLMs at cloud, leading to higher service delays. Meanwhile, Gemma 7B + Gemini 1.5 Pro can achieve the lowest overall delay than other combinations. This is because of the low TTFT values of these two LLMs, which are 1/155 second for Gemma 7B and 1/58 second for Gemini 1.5 Pro. Other LLMs have much higher values, i.e., 1/89 second for Llama2-7B and 1/40 second for Llama2-70B[11]. In addition, some supplementary experiment results and analyses can be found in [12].
@@ V Conclusion
GAI is a promising technique for future 6G networks, and this work investigates the generation task offloading problems in 6G edge-cloud. It proposes a novel in-context learning method for generation task offloading, and the simulations demonstrated that the proposed technique can achieve satisfactory generation service quality. It is worth noting that such an edge-cloud deployment method may also increase the vulnerabilities of LLMs, e.g., inference and extraction attacks. In the future, we will explore secure and robust deployment strategies for LLMs in 6G networks, better guaranteeing the generation service reliability and quality.